If f(1)=9 and f(n)=f(n-1)+2 then find the value of f(6)
Found 2 solutions by ikleyn, ewatrrr:
Answer by ikleyn(52776)

.
If f(1)=9 and f(n)=f(n-1)+2 then find the value of f(6)
~~~~~~~~~~~~~~~

The formula describes the arithmetic progression with the first term of 9
and the common difference of 2.


Therefore, f(6) is the 6th term of this progression

    f(6) = 9 + 5*2 = 9 + 10 = 19.    ANSWER

Good joking Math problem.



Answer by ewatrrr(24785)

If f(1)=9 and f(n)=f(n-1)+2 then find the value of f(6)
f(2) = 9 + 2 = 11
f(3) = 11 + 2 = 13
f(4) = 13 + 2 = 15
f(5) = 15 + 2 = 17
f(6) = 17 + 2 = 19
OR
as previously stated:
Arithmetic Series: d = 2
f(n) = 9 + (n-1)2